# Frequency Is What You Need: Word-frequency Masking Benefits Vision-Language Model Pre-training

### v2

@@ B.7 Applying text masking on SigLIP
Table 15: We apply text masking (75%) on SigLIP to pre-train on CC12M for 30 epochs with 75% image masking for speedup pre-training. Reporting the zero-shot classification accuracy on ImageNet-1K.
@@
Appendix C Ablation
Table 16: We pre-train CLIPF on CC12M dataset across different thresholds in Equation 2. The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
@@ Appendix C Ablation
C.1 Threshold Analysis
To investigate the impact of the threshold $t$ in Equation 2 on model performance, we pre-trained models with varying thresholds. As shown in Table 16, thresholds of $51e-$ 5, $61e-$ 6, and $71e-$ 7 yield comparable results, all outperforming the other text masking strategies. This indicates that CLIPF is relatively insensitive to small variations in thresholds, which are intended to maintain the word masking probability within the range of 0 to 1. However, using too small a threshold reduces the differences in text masking probabilities; therefore, we do not recommend using an excessively low threshold.
@@ C.1 Threshold Analysis
C.2 Word and Token Analysis
Since Open_CLIP encodes text using byte-pair encoding (BPE) [9], some words are represented by multiple tokens. Therefore, in this experiment, we calculate the masking probability based on token frequency rather than whole words. Table 17 compares the performance of both approaches. Although token frequency masking may disrupt the original word structure and thus achieve lower performance compared to word frequency masking, it still slightly outperforms other text masking strategies.

### v3

@@ C.6 Applying text masking on SigLIP
Table 18: Comparison of the ImageNet-1k zero-shot accuracy performance of SigLIP with different text masking strategies. The models are pre-trained on CC12M [3] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
Appendix D Ablation
@@
Table 19: We pre-train CLIPF on the CC12M dataset across different thresholds in Equation 2. The models are pre-trained on CC12M [3] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
@@
Model Masking Time (minutes)
FLIP trancation 39.81
random 48.43
CLIPA block 40.81
syntax 217.76
CLIPF frequency 84.64
Table 21: Processing times for the CC12M dataset by using different masking strategies.
D.1 Threshold Analysis
To investigate the impact of the threshold $t$ in Equation 2 on model performance, we pre-trained models with varying thresholds. As shown in Tab. 19, thresholds of 1e51e-5, 1e61e-6, and 1e71e-7 yield comparable results, all outperforming the other text masking strategies. This indicates that CLIPF is relatively insensitive to small variations in thresholds, which are intended to maintain the word masking probability within the range of 0 to 1. However, using too small a threshold reduces the differences in text masking probabilities; therefore, we do not recommend using an excessively low threshold.
D.2 Word and Token Analysis
Since Open_CLIP encodes text using byte-pair encoding (BPE) [15], some words are represented by multiple tokens. Therefore, in this experiment, we calculate the masking probability based on token frequency rather than whole words.  Tab. 20 compares the performance of both approaches. Although token frequency masking may disrupt the original word structure and thus achieve lower performance compared to word frequency masking, it still slightly outperforms other text masking strategies.
Appendix E Computational overhead
We carried out a simple measurement of how long it takes to tokenize CC12M using one thread. As shown in  Tab. 21, syntax masking required 2.5 times the number of minutes of CLIPF (218 vs. 85) and about 5.4 times that of truncation, random, and block masking.